suppose that 1 out of 10 plasma televisions shipped with a defective speaker. out of a shipment of n=400 plasma televisions. find the probability that there are
a) at most 40 with defective speakers (Hint: Use the dishonest- coin principle with P= 1/10=0.1 to the find the mean and standard deviation.
B) Most than 52 with defective speakers
1
Expert's answer
2019-04-24T09:59:18-0400

Dishonest-Coin Principle

Let X denote the number of heads in n tosses of a coin (assume "n\geq30)". Let p denote the probability of heads on each toss.Then, X has an approximately normal distribution with mean "\mu=n\cdot p" and standard deviation "\sigma =\sqrt{n\cdot p\cdot (1-p)}".

We have that "p=0.1, n=400". Then "\mu=400\cdot 0.1=40, \sigma =\sqrt{400\cdot 0.1\cdot (1-0.1)}=6."

a) at most 40 with defective speakers


"z={x-\mu \over \sigma}={40.5-40 \over 6}=0.083333"

"P(X\leq40)=P(X<40.5)=P(Z<0.083333)=0.533207"

b) Most than 52 with defective speakers


"z={x-\mu \over \sigma}={52.5-40 \over 6}=2.083333"

"P(X>52)=P(X>52.5)=0.018610"